• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Sulkosky

    Log entry time 13:05:14 on July21,2003

    Entry number 105841

    keyword=Transverse Laser Settings

    JP, Jing, and I adjusted the Transverse laser temperatures to adjust the laser wavelengths.

    Laser #1 Temperature changed from 7.2C to 8.5C
    Laser #2 Temperature changed from 10.8C to 11.4C
    Laser #3 Temperature changed from 15.3C to 14.8C

    Note: The Transverse laser absorption spectrum can be seen entirely with one or two lasers on.